OpenStudy (anonymous):

Let f be the function defined by f(x) = x-2cosx on the closed interval [0,2pi] Determine the value of x at which f has its -Absolute Max -Absolute Min

OpenStudy (anonymous):

To find the absolute min and max, I know I have to take the derivative I got: 1+2sinx = 0 sinx = -1/2 Now is this an arcsin problem? x = arcsin(-1/2) ? ? Where do I go from there?

OpenStudy (anonymous):

I'm still here :)

OpenStudy (anonymous):

the derivative is positive for every x which belongs to [0,2pi]. Means, the function is monotonically increasing. From that, you just find f(0) and f(2pi), and these are the min and the max..

OpenStudy (anonymous):

sinx=-0.5 only for the angle -30 degrees (-pi/6), which is not between 0 to 2pi.

OpenStudy (anonymous):

So I can make the derivatives equal to 0 and 2pi to find min and max?

OpenStudy (anonymous):

No. You found out with the derivative that the function is monotonically increasing. So it's minimum is the left value=0, and the highest point of it's graph is the right value (2pi). To find these values, you put x=0 and x=2pi in the function, not in the derivative.

OpenStudy (anonymous):

f(0)=0-2cos(0)=-2 f(2pi)=2pi-2cos(2pi)=2pi-2

OpenStudy (anonymous):

Minimum: (0,-2) ; Maximum: (2pi,2-2pi)

OpenStudy (anonymous):

Oh the only reason why I found the derivative to see if it was increasing or decreasing. OH that's how I got those values. OKay thank you :)

OpenStudy (anonymous):

:D